OpenStudy (anonymous):

A gardener is planning to plant a rectangular vegetable garden during the spring. The rectangle below represents a plan for the garden with a width of x feet. Find the length of the rectangle if the area of a rectangle is x^2 +12x. (hint: factor the trinomial x^2 +12x +0 using the guess and check method.)

OpenStudy (anonymous):



zepdrix (zepdrix):

Recall that Area of a rectangle is \(\large\rm A=\ell\cdot w\)

zepdrix (zepdrix):

So we have:\[\large\rm x^2+12x=x\cdot\ell\]Ya?

zepdrix (zepdrix):

Try to factor something out of the left side of the equation. Do they have anything in common?

OpenStudy (anonymous):

is that all i have to do for a?

zepdrix (zepdrix):

I dunno :P I don't wanna download a document lol a picture would be better +_+

OpenStudy (anonymous):

-.-